Find and classify the differential equation by eliminating a,b and c from general solution equation : y=aex+be2x+ce-3+x2

where a ,b ,and c are arbitrary constants
Answer #1

Given. y = ae tbe tee ²x2 N differentiating with respect ton, y = al + 2be+21-0 y= ae+4be?4+2 (again differentiating wiIf the Given equation is, y =ae +be+cē3*+x2 y = ae* + 2be?- 3cē3*+ en Y = a + 4be² +9ce342 Y = ae +8 bez _ 27 00 34 Re R